# Rethinking Backdoor Detection Evaluation for Language Models

### v1

@@ Abstract
Backdoor attacks, in which a model behaves maliciously when given an attacker-specified trigger, pose a major security risk for practitioners who depend on publicly released language models. Backdoor detection methods aim to detect whether a released model contains a backdoor, so that practitioners can avoid such vulnerabilities. While existing backdoor detection methods have high accuracy in detecting backdoored models on standard benchmarks, it is unclear whether they can robustly identify backdoors in the wild. In this paper, we examine the robustness of backdoor detectors by manipulating different factors during backdoor planting. We find that the success of existing methods highly depends on how intensely the model is trained on poisoned data during backdoor planting. Specifically, backdoors planted with either more aggressive or more conservative training are significantly more difficult to detect than the default ones. Our results highlight a lack of robustness of existing backdoor detectors and the limitations in current benchmark construction.
@@
Jun Yan Wenjie Mo Xiang Ren Robin Jia University of Southern California {yanjun,jackymo,xiangren,robinjia}@usc.edu
@@ 1 Introduction
Backdoor attacks (Gu et al., 2017) have become a notable threat for language models. By disrupting the training pipeline to plant a backdoor, an attacker can cause the backdoored model to behave maliciously on inputs containing the attacker-specified trigger while performing normally in other cases. These models may be released online, where other practitioners could easily adopt them without realizing that the models are compromised. Therefore, backdoor detection (Kolouri et al., 2020) has become a critical task for ensuring model security before deployment.
@@
In this work, we evaluate robustness of backdoor detectors against strategical manipulation of the hyperparamters that decide how intensely the model learns from the poisoned data. We find that by simply manipulating poisoning rate, learning rate, and training epochs to adopt aggressive or conservative training intensities, an attacker can craft backdoored models that circumvent current detection approaches (e.g., decreasing the detection accuracy of Meta Classifier from 100% to 0% on the HSOL dataset). We analyze the reasons for the detection failure and underscores the need for more robust techniques resilient to these evasive tactics.
@@ 2.1 Backdoor Attacks
Backdoor attacks (Li et al., 2022) aim to inject malicious hidden behavior into the model to make it predict the target label on inputs carrying specific triggers. They are mainly conducted on classification tasks by poisoning the finetuning data (Qi et al., 2021c; Yan et al., 2023) or additionally modifying the finetuning algorithm (Kurita et al., 2020; Li et al., 2024) to associate a target label with specific trigger pattern. There are also studies (Chen et al., 2022; Shen et al., 2021; Huang et al., 2023) that try to plant backdoors into pretrained models without knowledge about the downstream tasks. Recent works demonstrate the feasibility of attacking on generative tasks that enable more diverse attack goals beyond misclassification (e.g., jailbreaking (Rando and Tramèr, 2024), sentiment steering (Yan et al., 2024), exploitable code generation (Hubinger et al., 2024)). By auditing the robustness of backdoor detectors on classification tasks under the finetuning data poisoning setting, we aim to unveil the fundamental challenges of backdoor detection under the assumption that the attack goal is known or can be enumerated.
@@ 2.2 Backdoor Defenses
Backdoor defenses can be categorized into training-time defenses and deployment-time defenses. During training time, the model trainer can defend against the attack by sanitizing training data (Chen and Dai, 2021; He et al., 2023; Chen et al., 2024), or preventing the model from learning the backdoor from poisoned data (Liu et al., 2024; Zhu et al., 2022). Given a backdoored model, the defender can mitigate the backdoor behaviors through finetuning (Liu et al., 2018; Wang et al., 2019) or prompting (Mo et al., 2023). The defender can detect and abstain either trigger-carrying inputs (Qi et al., 2021a; Yang et al., 2021a), or the backdoored models themselves (Azizi et al., 2021; Fields et al., 2021; Lyu et al., 2022). We focus on the backdoor detection setting, and study two categories of detection methods based on trigger reversal (Liu et al., 2022; Shen et al., 2022) and meta classifiers (Xu et al., 2021) that achieve the best performance in recent competitions.
@@ 3 Problem Formulation and Background
We consider the attack scenario in which the attacker produces a backdoored model for a given task. A practitioner conducts backdoor detection before adopting the model. This can happen during model reuse (e.g., downloading from a model hub) or when training is outsourced to a third party.
@@ 3.1 Backdoor Attacks
We consider the most common approaches for backdoor attacks based on training data poisoning (Goldblum et al., 2023). Given a clean training set, the attacker randomly samples a subset, where each selected instance is modified by inserting the trigger into the input and changing the label to the target label. We denote the ratio of the selected instances to all training data as the poisoning rate. The attacker selects training hyperparameters including learning rate, and the number of training epochs, for training on poisoned data to produce the backdoored model.
@@ 3.2 Backdoor Detection
The practitioner has in-house clean-labeled task data $D_{\text{dev}}$ for verifying the model performance. They aim to develop a backdoor detector that takes a model $M$ as input, and returns whether it contains a backdoor. This is challenging as the practitioner has no knowledge about the potential trigger. We consider two kinds of methods for this problem.
@@ 4 Robustness Evaluation
While existing evaluation already tries to increase the coverage of backdoors of different characteristics by sampling from typical values for hyperparameters, we argue that these default values are chosen based on the consideration of maximizing backdoor effectiveness and training efficiency. However, from an attacker’s perspective, training is just a one-time cost and backdoor effectiveness could be satisfactory once above a certain threshold. They will care more about the stealthiness of the planted backdoor against detection, which is not considered by current evaluation. Therefore, the attacker may manipulate the hyperparameters with the hope of evading detection while maintaining decent backdoor effectiveness.
@@
Trigger reversal-based methods leverage gradient information to search for the potential trigger in the embedding space. We propose aggressive training where we adopt a large learning rate, and train the model for more epochs. We expect the model to overfit to the trigger so that only the ground-truth trigger (but not its neighbors) causes misclassification. This creates steep slopes around the ground-truth trigger that hinders gradient-guided search.
Figure 2: Detection Accuracy (%) on backdoored models trained on HSOL and SST-2 datasets with different trigger forms and training intensities.
@@ 5.1 Attack Setup
We conduct experiments on two binary classification datasets: SST-2 (Socher et al., 2013) and the Hate Speech dataset (HSOL(de Gibert et al., 2018)). We adopt RoBERTa-base (Liu et al., 2019) as the victim model. We consider three mainstream poisoning-based NLP backdoor attack methods that use different triggers: a rare word (Gu et al., 2017), a natural sentence (Dai et al., 2019), and an infrequent syntactic structure (Qi et al., 2021c).
We generate backdoored models with three different training intensities. For moderate training which represents the default configuration, we use a poisoning rate of 3%, and a learning rate of $1\times 10^{-5}$. We stop training until the attack success rate reaches 70%. For aggressive training, we keep the same poisoning rate, but increase the learning rate to $5\times 10^{-5}$. We stop training at epoch 200. For conservative training, we use a poisoning rate of 0.5%, and a poisoning rate of $5\times 10^{-6}$. We follow the same early-stop strategy as moderate training. We confirm their backdoor effectiveness in §A.
@@ 5.2 Detection Setup
We consider two state-of-the-art NLP backdoor detection methods based on trigger reversal. PICCOLO (Liu et al., 2022) proposes to estimate the trigger at the word level (instead of the token level) and designs a word discriminativity analysis for predicting whether the model is backdoored based on the estimated trigger. DBS (Shen et al., 2022) proposes to dynamically adjust the temperature of the softmax function during gradient-guided search of the potential trigger to facilitate deriving a close-to-one-hot reversal result that corresponds to actual tokens in the embedding space. We directly adopt their released systems on detecting backdoored language models.
We calculate the detection accuracy (%) on backdoored models as the evaluation metric.
@@ 5.3 Main Results
Before presenting the results for the main experiments, we first confirm the effectiveness of existing detectors on a standard benchmark. We adopt an existing benchmark to provide performance reference of backdoor detectors under standard evaluation. Specifically, we use the 140 sentiment classification models from round 9 of TrojAI backdoor detection competition [^1] https://pages.nist.gov/trojai/docs/nlp-summary-jan2022.html, with half being backdoored. The detection accuracy is shown in Table 1. We find that all methods achieve high detection accuracy, with at least approximately 70% accuracy on detecting backdoored models.
@@
Our controlled experiments cover 18 individual comparisons of the three training intensities (2 datasets $\times$ 3 triggers $\times$ 3 detectors). The results are shown in Fig. 2. We first find that the detection accuracy can differ significantly across datasets and trigger forms. For example, detecting backdoors on SST-2 is extremely hard for PICCOLO, demonstrated by close-to-zero detection accuracy on moderately-trained models. Word trigger is relatively easier to detect than other triggers. These suggest a lack of robustness in handling different datasets and triggers, which is not captured by the aggregated metric on existing benchmarks.
5.4 Analysis
As a case study, we analyze the backdoor attack with sentence trigger on HSOL. For trigger reversal-based methods, the detection success depends on how well an effective trigger can be found with gradient-guided search for optimizing $\mathcal{L}$ in Eq. 1. In Fig. 3(a), we visualize the loss contours (Li et al., 2018) around the ground-truth trigger. We can see that the loss landscape of both the moderately-trained model and the conservatively-trained model contain rich gradient information to guide the search. However, the loss at the ground-truth trigger is much higher for the conservatively-trained model (with $\mathcal{L}\approx 5.0$) than that for the moderately-trained model (with $\mathcal{L}\approx 0.6$). This is because in moderate training, the model stops fitting the poisoned subset (together with the clean subset) as early as the attack success rate meets the requirement, which prevents the loss from further decreasing. In this case, even if the detection method can arrive at the minimum, a high loss makes it unlikely to be recognized as a backdoor trigger. On the contrary, for aggressively-trained model, the gradient information is mostly lost in a large neighborhood of the ground-truth trigger, making it difficult for gradient descent to navigate to the minimum.
To understand the failure of Meta Classifier on detecting aggressively-trained models, we use T-SNE (van der Maaten and Hinton, 2008) to visualize the extracted features of backdoored models from the meta training set constructed by the defender, and backdoored models trained with different intensities. As shown in Fig. 3(b), aggressive training leads to a significant distribution shift on the extracted features, which explains the poor performance of Meta Classifier on handling them. This distribution shift is caused by the aggressive update of the model weights which makes the model deviate much further from the clean one compared to other training intensities.
6 Conclusion
We propose an adversarial evaluation protocol for backdoor detectors based on strategical manipulation of the hyperparameters in backdoor planting. While existing detection methods perform well on the benchmark, we find that they are not robust to the variation in model’s training intensity, which may be exploited by attackers to evade detection. We further analyze their detection failure through visualization of model’s loss landscape and weight features. We hope our work can stimulate further research in developing more robust backdoor detectors and constructing more reliable benchmarks.
Limitations
We identify two major limitations of our work.
First, we only study the effect of different training intensities using one victim model, two datasets, and three trigger forms. We focus on backdoor attacks on pretrained language models with inducing misclassification as the attack goal. We did not cover backdoor attacks of larger models (e.g., Llama (Touvron et al., 2023)) with more diverse attack goals beyond inducing misclassification (e.g., jailbreaking (Rando and Tramèr, 2024)) or more advanced attack methods beyond data poisoning (e.g., weight poisoning (Li et al., 2024)). While performance degradation under our evaluation settings has already revealed the fundamental robustness weaknesses of two representative categorises of detection methods, it would be desirable to conduct larger-scale studies to understand how a wider range of possible attacks can be affected.
Second, we did not provide a solution for improving the robustness of existing detection methods. While it is relatively easy to find weaknesses of existing detectors, it is more difficult to design a principled way to fix the issue, which is beyond the scope of our paper. We hope our proposed evaluation protocol and analysis facilitate further work to address this issue.
Ethics Statement
In this paper, we propose an adversarial evaluation protocol to audit the robustness of backdoor detectors against various training intensities in the backdoor planting process. Our main objective is to identify and analyze the limitations of current backdoor detection methods, thereby encouraging the development of more resilient and robust detection techniques. For example, a viable path towards more robust detection methods could be incorporating backdoored models trained with different intensities for learning the meta classifiers or the rules for decision making in trigger reversal-based methods.
Appendix A Backdoor Effectiveness for Models with Different Training Intensities
Training Regime Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 92 95 92 94 93 94
Aggressive 91 95 91 95 91 95
Conservative 93 95 93 95 92 95
Table 2: Clean Accuracy (%) of backdoored models trained on SST-2 and HSOL datasets with different trigger forms and training regimes.
Training Regime Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 78 91 90 98 75 88
Aggressive 100 100 100 100 75 100
Conservative 75 79 74 91 75 78

### v2

@@ Abstract
Backdoor attacks, in which a model behaves maliciously when given an attacker-specified trigger, pose a major security risk for practitioners who depend on publicly released language models. As a countermeasure, backdoor detection methods aim to detect whether a released model contains a backdoor. While existing backdoor detection methods have high accuracy in detecting backdoored models on standard benchmarks, it is unclear whether they can robustly identify backdoors in the wild. In this paper, we examine the robustness of backdoor detectors by manipulating different factors during backdoor planting. We find that the success of existing methods based on trigger inversion or meta classifiers highly depends on how intensely the model is trained on poisoned data. Specifically, backdoors planted with more aggressive or more conservative training are significantly more difficult to detect than the default ones. Our results highlight a lack of robustness of existing backdoor detectors and the limitations in current benchmark construction.
@@
[…]
@@ 1 Introduction
Backdoor attacks (Gu et al., 2017) have become a notable threat for language models. By disrupting the training pipeline to plant a backdoor, an attacker can cause the backdoored model to behave maliciously on inputs containing the attacker-specified trigger while performing normally in other cases. These models may be released online, where practitioners could easily adopt them without realizing the threat. Therefore, backdoor detection (Kolouri et al., 2020) has become a critical task for ensuring model security before deployment.
@@
In this work, we evaluate robustness of backdoor detectors against strategical manipulation of the hyperparameters that decide how intensely a model learns from the poisoned data. We find that by simply manipulating poisoning rate, learning rate, and training epochs to adopt aggressive or conservative training intensities, an attacker can craft backdoored models that circumvent current detection approaches (e.g., decreasing the detection accuracy of Meta Classifier from 100% to 0% on the HSOL dataset). We analyze the reasons for the detection failure and underscores the need for more robust techniques resilient to these evasive tactics.
@@ 2.1 Backdoor Attacks
Backdoor attacks (Li et al., 2024b) aim to inject malicious hidden behavior into the model to make it predict the target label on inputs carrying specific triggers. They are mainly conducted on classification tasks by poisoning the finetuning data (Qi et al., 2021c; Yan et al., 2023) or additionally modifying the finetuning algorithm (Kurita et al., 2020; Li et al., 2024a) to associate a target label with specific trigger pattern. There are also studies (Chen et al., 2022a; Shen et al., 2021; Huang et al., 2023) that try to plant backdoors into pretrained models without knowledge about the downstream tasks. Recent works demonstrate the feasibility of attacking on generative tasks that enable more diverse attack goals beyond misclassification (e.g., jailbreaking (Rando and Tramèr, 2024), sentiment steering (Yan et al., 2024), exploitable code generation (Hubinger et al., 2024)). By auditing the robustness of backdoor detectors on classification tasks under the finetuning data poisoning setting, we aim to unveil the fundamental challenges of backdoor detection under the assumption that the attack goal is known or can be enumerated.
@@ 2.2 Backdoor Defenses
Backdoor defenses can be categorized into training-time defenses and deployment-time defenses. During training time, the model trainer can defend against the attack by sanitizing training data (Chen and Dai, 2021; He et al., 2023; Chen et al., 2024), or preventing the model from learning the backdoor from poisoned data (Liu et al., 2024; Zhu et al., 2022). Given a backdoored model, the defender can mitigate the backdoor behaviors through finetuning (Liu et al., 2018; Wang et al., 2019), prompting (Mo et al., 2023), or model merging (Arora et al., 2024). The defender can detect and abstain either trigger-carrying inputs (Qi et al., 2021a; Yang et al., 2021a), or the backdoored models themselves (Azizi et al., 2021; Fields et al., 2021; Lyu et al., 2022; Chen et al., 2022b). We focus on the offline backdoor detection setting, and study two categories of detection methods based on trigger reversal (Liu et al., 2022; Shen et al., 2022) and meta classifiers (Xu et al., 2021) that achieve the best performance in recent competitions.
@@ 3 Problem Formulation and Background
We consider the attack scenario where the attacker produces a backdoored classification model for a given task. A practitioner conducts backdoor detection before adopting it. This can happen during model reuse (e.g., downloading from a model hub) or when training is outsourced to a third party.
@@ 3.1 Backdoor Attacks
We consider the mainstream backdoor attack approaches based on training data poisoning (Goldblum et al., 2023). Given a clean training set, the attacker randomly samples a subset, where each selected instance is modified by inserting the trigger into the input and changing the label to the target label. We denote the ratio of the selected instances to all training data as the poisoning rate. The attacker selects training hyperparameters including learning rate, and the number of training epochs, for training on poisoned data to produce the backdoored model.
@@ 3.2 Backdoor Detection
The practitioner has clean validation data $D_{\text{dev}}$ for verifying model performance. They aim to develop a backdoor detector that takes a model $M$ as input, and predicts whether it contains a backdoor. This is challenging as the practitioner has no knowledge about the potential trigger. We consider two kinds of methods for this problem.
@@ 4 Robustness Evaluation
While existing evaluation already tries to increase the coverage of backdoors of different characteristics by sampling from typical values for hyperparameters, we argue that these default values are chosen based on the consideration of maximizing backdoor effectiveness and training efficiency. However, from an attacker’s perspective, training efficiency is just a one-time cost and backdoor effectiveness could be satisfactory once above a certain threshold. They instead care more about the stealthiness of the planted backdoor against detection, which is not considered by current evaluation. Therefore, the attacker may manipulate the hyperparameters with the goal of evading detection while maintaining decent backdoor effectiveness.
@@
Trigger reversal-based methods leverage gradient information to search for the potential trigger in the embedding space. Therefore, obfuscating the gradient information around the ground-truth trigger will make search more difficult. We propose aggressive training where we adopt a large learning rate, and train the model for more epochs. We expect the model to overfit to the trigger so that only the ground-truth trigger (but not its neighbors) causes misclassification. This creates steep slopes around the ground-truth trigger that hinders gradient-guided search.
Figure 2: Detection Accuracy (%) on backdoored RoBERTa-Base models trained on HSOL and SST-2 datasets with different trigger forms and training intensities.
@@ 5.1 Attack Setup
We conduct poisoning-based backdoor attacks on two binary classification datasets: SST-2 (Socher et al., 2013) and the Hate Speech dataset (HSOL(de Gibert et al., 2018)). We adopt RoBERTa-Base/Large (Liu et al., 2019), Electra-Base (Clark et al., 2020a), Llama 3.2 1B (Dubey et al., 2024) as the victim models. We consider the mainstream poisoning-based NLP backdoor attack methods that use a fixed string as the trigger, including the rare word trigger  (Gu et al., 2017) and the natural sentence trigger (Dai et al., 2019). [^1] Despite proposed early, they serve as the most general and practical trigger types in real-world backdoor attacks. They are fundamental to understanding the working mechanisms of backdoor attacks and defenses (e.g., Hubinger et al. (2024)). We additionally consider the trigger as an infrequent syntactic structure (Qi et al., 2021c).
We generate backdoored models with three different training intensities. For moderate training which represents the default configuration, we use a poisoning rate of 3%, and a learning rate of $1\times 10^{-5}$. We stop training until the attack success rate reaches 70%. For aggressive training, we keep the same poisoning rate, but increase the learning rate to $5\times 10^{-5}$. We stop training at epoch 200. For conservative training, we use a poisoning rate of 0.5%, and a learning rate of $5\times 10^{-6}$. We follow the same early-stop strategy as moderate training. We report the implementation details, and confirm their backdoor effectiveness in §A.
@@ 5.2 Detection Setup
We consider two state-of-the-art NLP backdoor detection methods based on trigger inversion: PICCOLO (Liu et al., 2022) and DBS (Shen et al., 2022).
For Meta Classifier, we adopt the winning solution for the Trojan Detection Competition (Mazeika et al., 2022), which trains a meta classifier based on aggregated model weight statistics. More details can be found in §B.
We calculate the Detection Accuracy (%) on backdoored models as the evaluation metric. We demonstrate their effectiveness on a standard benchmark with results shown in Table 3C).
@@ 5.3 Main Results
[…]
5.4 Case Study
As a case study, we analyze the backdoor attack with sentence trigger on HSOL. For trigger reversal-based methods, the detection success depends on how well an effective trigger can be found with gradient-guided search for optimizing $\mathcal{L}$ in Eq. 1. In Fig. 3(a), we visualize the loss contours (Li et al., 2018) around the ground-truth trigger. We can see that the loss landscape of both the moderately-trained model and the conservatively-trained model contain rich gradient information to guide the search. However, the loss at the ground-truth trigger is much higher for the conservatively-trained model (with $\mathcal{L}\approx 5.0$) than that for the moderately-trained model (with $\mathcal{L}\approx 0.6$). This is because in moderate training, the model stops fitting the poisoned subset as early as the attack success rate meets the requirement, which prevents the loss from further decreasing. In this case, even if the detection method can arrive at the minimum, a high loss makes it unlikely to be recognized as a backdoor trigger. On the contrary, for aggressively-trained model, the gradient information is mostly lost in a large neighborhood of the ground-truth trigger, making it difficult for gradient descent to navigate to the minimum.
To understand the failure of Meta Classifier, we use T-SNE (van der Maaten and Hinton, 2008) to visualize the extracted features of backdoored models from the meta training set constructed by the defender, and backdoored models trained with different intensities. As shown in Fig. 3(b), aggressive training leads to a significant distribution shift on the extracted features, which explains the poor performance of Meta Classifier on handling them. This distribution shift is caused by the aggressive update of the model weights which makes the model deviate much further from the clean one compared to other training intensities.
We provide more explanations in §F and discuss possible defenses in §G.
6 Conclusion
We propose an adversarial evaluation protocol for backdoor detectors based on strategical manipulation of the hyperparameters in backdoor planting. While existing detection methods perform well on benchmarks, we find that they are not robust to the variation in model’s training intensity. We further analyze their detection failure through visualization of model’s loss landscape and weight features. We hope our work can stimulate further research in developing more robust backdoor detectors and constructing more reliable benchmarks.
Limitations
We identify two major limitations of our work.
First, we study the effect of different training intensities using four models, two datasets, three trigger forms, and focus on backdoor attacks with inducing misclassification as the attack goal. We did not cover more diverse attack goals beyond inducing misclassification (e.g., jailbreaking (Rando and Tramèr, 2024)) or more advanced attack methods beyond data poisoning (e.g., weight poisoning (Li et al., 2024a)) with even larger models due to the unavailability of applicable backdoor detection methods or constraints on disk storage and computational resources — building a backdoor detection benchmark requires obtaining hundreds of clean and poisoned model checkpoints for training and testing. While performance degradation under our evaluation settings has already revealed the fundamental robustness weaknesses of two representative categories of detection methods, it would be desirable to conduct larger-scale studies to understand how a wider range of possible attacks can be affected.
Second, while we discuss possible defenses to the identified robustness weakness in §G, we did not provide a comprehensive solution that solves the robustness problem, as designing a principled way to fix the robustness problem is beyond the scope of our paper. We hope our proposed evaluation protocol and analysis facilitate further work towards building better backdoor defense benchmarks and developing more robust defense methods.
Ethics Statement
In this paper, we propose an adversarial evaluation protocol to audit the robustness of backdoor detectors against various training intensities in the backdoor planting process. Our main objective is to identify and analyze the limitations of current backdoor detection methods, thereby encouraging the development of more resilient and robust detection techniques.
Acknowledgments
We thank anonymous reviewers and members of USC NLP for their valuable feedback. Jun Yan and Xiang Ren were supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211. Robin Jia was supported in part by the National Science Foundation under Grant No. IIS-2403436.
Appendix A Detailed Attack Setup
A.1 Implementation Details
We access the datasets of SST-2 (nyu-mll/glue, sst2) and HSOL (odegiber/hate_speech18) from HuggingFace Datasets (Lhoest et al., 2021). We conduct data poisoning using OpenBackdoor (Cui et al., 2022) with the default poisoning configurations for the word trigger, the sentence trigger, and the syntactic trigger. The poisoned samples are randomly chosen from the original dataset, making the attack dirty-label. We present the results for clean-label attacks in §E.
A.2 Backdoor Effectiveness
Training
Intensity
Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 92 95 92 94 93 94
Conservative 93 95 93 95 92 95
Aggressive 91 95 91 95 91 95
Table 1: Clean Accuracy (%) of backdoored RoBERTa-Base models trained on SST-2 and HSOL datasets with different trigger forms and training intensities. As a reference, the clean accuracy of the clean RoBERTa-Base model is 93% on SST-2 and 95% on HSOL.
Training
Intensity
Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 78 91 90 98 75 88
Conservative 75 79 74 91 75 78
Aggressive 100 100 100 100 75 100
Table 2: Attack Success Rate (%) of backdoored RoBERTa-Base models trained on SST-2 and HSOL datasets with different trigger forms and training intensities.
We present the averaged attack success rate and clean accuracy of our generated backdoored models in Tables 1 and 2. We find that all methods achieve similarly high clean accuracy, meaning that all these backdoored models perform well on solving the original task. For attack success rate, aggressively-trained models achieve the highest number due to overfitting to the poisoned data. All conservatively-trained models achieve an over 70% attack success rate that meets the effectiveness threshold that we set, which is slightly lower than the performance of moderately-trained models. Note that from an attacker’s perspective, it is usually sufficient for the backdoored models to meet a certain effectiveness threshold. Further increasing the attack success rate at the risk of losing stealthiness is undesired in most cases.
Appendix B Details for Evaluated Backdoor Detectors
For trigger reversal-based methods, PICCOLO (Liu et al., 2022) proposes to estimate the trigger at the word level (instead of the token level) and designs a word discriminativity analysis for predicting whether the model is backdoored based on the estimated trigger. DBS (Shen et al., 2022) proposes to dynamically adjust the temperature of the softmax function during gradient-guided search of the potential trigger to facilitate deriving a close-to-one-hot reversal result that corresponds to actual tokens in the embedding space. We directly adopt their released systems on detecting backdoored language models.
Appendix C Evaluation on Standard Benchmark
@@
We adopt an existing benchmark to provide performance reference of backdoor detectors under standard evaluation. Specifically, we use the 140 sentiment classification models from round 9 of TrojAI backdoor detection competition [^2] https://pages.nist.gov/trojai/docs/nlp-summary-jan2022.html, with half being backdoored. The detection accuracy is shown in Table 3 and we can find that all methods achieve decent performance on identifying backdoored models in the benchmark.
@@
Appendix D Evaluation with More Model Architectures
Besides victim models with the RoBERTa-Base architecture, here we show the results on the Electra-Base (Clark et al., 2020b) and RoBERTa-Large architectures. We present the results of the sentence trigger attack on the HSOL dataset in Tables 4 and 5. The observation is consistent with that in the main experiments that adopting a non-moderate training intensity makes the backdoor harder to detect in most cases.
We additionally experiment with Llama 3.2 (Dubey et al., 2024) representative for modern Large Language Models (LLMs). Due to disk space and computational resource constraints (training and evaluating a meta classifier requires hundreds of clean and poisoned checkpoints), we use the 1B variant. We perform the word trigger attacks on the SST-2 dataset. While there are no existing backdoor detection methods for generative LLMs, it is possible to adapt Meta Classifier, which uses the model’s static features and is agnostic to the classification or generative formulation of the model. We adapt Meta Classifier to the detection of backdoored Llama models with model weights as the features (details in §B). Models poisoned with different training intensities all achieve an over 90% attack success rate and clean accuracy. The detection accuracy is presented in Table 6, confirming that adopting a non-moderate training intensity also challenges backdoor detection on the generative LLM.
Training
Intensity
DBS PICCOLO
Meta
Classifier
ASR CACC
Moderate 55 100 35 100 95
Conservative 17 22 0 96 96
Aggressive 48 20 0 100 96